Am doing this right? Is the ordered pair 5,-4 the answer for the following question
which ordered pair is a solution to 3x - 4y = -1
thanks for your help

Found 2 solutions by stanbon, ice_water35:
Answer by stanbon(75887)

Substituting get 3(5)-4(-4)= 15+16=31
Cheers,
Stan H.

Answer by ice_water35(8)

>>>>>5 and -4 is not the answer. Here it goes:
3x - 4y = -1
3(5)-4(-4)=-1
15-(-16)=-1
15+16=1
31=1 >this is incorrect
if you want to know the correct answer, just substitute the given ordered pair to the given equation. And if it fits the equation, that ordered pair is correct.
If you have further questions, you may contact me. This is my number: 09223011043